OpenStudy (anonymous):

Which does not represent y as a function of x? A. x=y^2+2 B. y+x^2+2 C. x=y+8 D. y=-x+8 I've been out of math for 3 years. Please help! :)

OpenStudy (anonymous):

anyone??

OpenStudy (anonymous):

did you mean to write B without an equal sign in it? If so, then I believe B is not a function

OpenStudy (anonymous):

I mean y as a function of x

OpenStudy (anonymous):

oh no B is y=x^2+2 sorry

OpenStudy (anonymous):

could you tell me how you figured that out please??

OpenStudy (anonymous):

my answer is wrong. Let me look at it again. One moment

OpenStudy (anonymous):

kay

OpenStudy (anonymous):

A. x=y^2+2 B. y=x^2+2 C. x=y+8 ------> can be written as y=x-8 (brought 8 over to the other side) D. y=-x+8 ok so now we have A. x=y^2+2 B. y=x^2+2 C. y=x-8 D. y=-x+8 they're all y as a function of x except for 1

OpenStudy (anonymous):

So A. is written "x in terms of y" because here \[x=y^2+2\]

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

@radar can explain this better than I can. He should be here momentarily

OpenStudy (anonymous):

A. y=sqrt(x-2) or y=-sqrt(x-2) It depends on whether y is positive or negative

OpenStudy (radar):

If a value of x results in more than 1 value of y, then it is not a function.

OpenStudy (anonymous):

ok that makes sense

OpenStudy (radar):

Looking at A. x=y^2 +2, then y^2=x-2 then \[y=sqrt{x-\2\}\\]Letting x=6, then\[y=sqrt{\6\-\2\}\=sqrt{\4\}\=\ \+\2\ and \-\2\\]A. is not a function.

OpenStudy (radar):

Looks like the equation editor is on the blink!

OpenStudy (anonymous):

could you guys help with one more problem??

OpenStudy (anonymous):

sure post it as a new problem on the left